MODS1 Commissioning Run 5: UTC 2011 February 19

Observer(s): R. Pogge, J. Antognini, & P. Martini (OSU)
Telescope Operator: D. Gonzalez-Huerta (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: J. Urban (LBTO)
Instrument Support: J. Morris (LBTO)

Summary

Preparations

Dewars filled at 4pm, all still holding >24h.

Worked on some small updates to the MODS control panel to implement the new instrument configuration commands.

Details

All times are UTC unless otherwise noted.

0100 - Overcast, winds 10-15m/s with gusts up to 18m/s. Temp -1°C, Humidity 80%. A storm system is on its way into the area, and there are rain bands out in the valley towards Willcox. Not looking good for tonight.

0500 - Overcast, 100% Humidity, Winds 15m/s from the SW gusting to near 20m/s. Temp -1°C, but no precipitation at this time.

0800 - Enveloped in fog but no precipitation as yet. Temp still -1°C, wind 10-15m/s from the SW with gusts to 18m/s.

1130 - Overcast, in and out of fog, but no precipitation yet. Wind and humidity mostly unchanged. We're calling it a night

-- RichardPogge - 19 Feb 2011
Topic revision: r4 - 19 Feb 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback